### v10

# Generalized domination structure in cubic graphs

Abstract
The minimum dominating set problem asks for a dominating set with minimum size. First, we determine some vertices contained in the minimum dominating set of a graph. Applying a particular scheme to the graph, the resulting graph is 2-connected and the length of each formed induced cycle is 0 mod 3. Label every three vertices for the 0 mod 3 induced cycles. Then there is a way of labeling in which the set of all labeled vertices is the minimum dominating set of the resulting graph, and is contained in the minimum dominating set of the original graph. We also consider the remaining vertices of the minimum dominating set of the original graph and determine all vertices contained in the minimum dominating set of a cubic graph. The complexity of the minimum dominating set problem for cubic graphs was shown to be APX-complete in 2000 and this problem is solved by our arguments in polynomial time.
keywords : minimum dominating set, cubic graph, NP-complete
MSC : 05C69
1 Introduction
This paper takes the minimum dominating set problem and aims to determine the minimum dominating set of a cubic graph. A cubic graph is a kind of graph, with all vertices degree 3. A dominating set of a graph $G$ is a set $S$ of vertices of $G$ such that every vertex $v$ of $G$ is either in $S$ or adjacent to a vertex of $S$. A minimum dominating set is a dominating set with minimum size.
First, we determine some vertices contained in the minimum dominating set of a graph. Applying a particular scheme to the graph, the resulting graph is 2-connected and the length of each formed induced cycle is 0 mod 3. The scheme has been relaxed from the published paper [3]. Label every three vertices for the 0 mod 3 induced cycles. Then there is a way of labeling in which the set of all labeled vertices is the minimum dominating set of the resulting graph, and is contained in the minimum dominating set of the original graph. We also consider the remaining vertices of the minimum dominating set of the original graph and determine all vertices contained in the minimum dominating set of a cubic graph.
The complexity of the minimum dominating set problem for cubic graphs was shown to be APX-complete [1] and this problem is solved by our arguments in polynomial time, which implies that all NP-complete problems are contained in P, but does not claim that P = NP in general.
2 Notation
In this paper, a graph $G$ is finite, undirected, and simple with the vertex set $V$ and edge set $E$. We follow the notations presented in [2]. For a vertex $v\in V(G)$, the open neighborhood, denoted by $N_{G}(v)$, is $\{u\in V(G)\colon\ uv\in E(G)\}$, also for a set $W\subseteq V(G)$, let $N_{G}(W)=\bigcup_{v\in W}N_{G}(v)$ and $N_{G}[W]=N_{G}(W)\cup W$. A dominating set $X\subseteq V(G)$ is such that $N_{G}[X]=V(G)$. A minimum dominating set, called a d-set, is a dominating set with the minimum size. $X$-3-path is a path with labels every 3 vertices. Two cycles $C_{1}$ and $C_{2}$ are said to be connecting without seams if $X$-3-paths can be assigned to $C_{1}$ and $C_{2}$ as $C_{1}\cap C_{2}$ is one non-trivial $X$-3-path.
3 Main results
We consider a connected graph $G$, otherwise consider each component one by one. We introduce the construction scheme ${\bf K}$ as follows.
${\bf K}$: Input a connected graph $G$.
(1) Let $G_{0}=G$ and $k=0$.
(2) If there exists a cut vertex in $G_{k}$, then proceed as follows. Let $v$ be the cut vertex in $G_{k}$. For every pair of components $C_{1}$ and $C_{2}$ of $G_{k}-v$ and for every pair of vertices $v_{1}\in C_{1}\cap N_{G_{k}}(v)$ and $v_{2}\in C_{2}\cap N_{G_{k}}(v)$, add an edge $v_{1}v_{2}$ to $G_{k}$. Increase $k$ by $1$.
(3) Repeat (2) while they occur.
(4) Find an induced cycle in $G_{k}$.
(4-1) If it is an induced cycle of length 0 mod 3, assign $X$-3-path to it. Increase $k$ by $1$.
(4-2) If it is an induced cycle of length 2 mod 3, then proceed as follows. Let $D_{2}$ be the induced cycle of length 2 mod 3. Take a 4-path $Q\subseteq D_{2}$, and set $Q=w_{1}w_{2}w_{3}w_{4}$. Add three edges $w_{1}w_{3}$, $w_{2}w_{4}$, and $w_{1}w_{4}$ to $G_{k}$. Assign $X$-3-paths to the induced cycles connecting without seams in $D_{2}+w_{1}w_{3}+w_{2}w_{4}+w_{1}w_{4}$. Increase $k$ by $1$.
(4-3) If it is an induced cycle of length 1 mod 3, then proceed as follows. Let $D_{1}$ be the induced cycle of length 1 mod 3. Take a 4-path $Q\subseteq D_{1}$, and set $Q=w_{1}w_{2}w_{3}w_{4}$. Add two edges $w_{1}w_{3}$ and $w_{2}w_{4}$ to $G_{k}$. Assign $X$-3-paths to the induced cycles connecting without seams in $D_{1}+w_{1}w_{3}+w_{2}w_{4}$. Increase $k$ by $1$.
(5) Find a shortest path $P$ between two vertices on $X$-3-path in $G_{k}$ that $X$-3-path is not assigned to.
(5-1) If it forms an induced cycle of length 0 mod 3 with the previously assigned $X$-3-path in $G_{k}$, assign $X$-3-path to it. Increase $k$ by $1$.
(5-2) If it forms an induced cycle of length 2 mod 3 with the previously assigned $X$-3-path in $G_{k}$, then proceed as follows. Let $D_{2}$ be the induced cycle of length 2 mod 3. If $|V(P-D_{2})|=1$, then set $u\in V(P-D_{2})$ and add all edges $uv$ such that $v\in V(D_{2}-P)$ to $G_{k}$. Otherwise, take a 4-path $Q\subseteq D_{2}$ such that $|V(P\cap Q)|\geq 3$, and set $Q=w_{1}w_{2}w_{3}w_{4}$. Add three edges $w_{1}w_{3}$, $w_{2}w_{4}$, and $w_{1}w_{4}$ to $G_{k}$. Assign $X$-3-paths to the induced cycles connecting without seams in $D_{2}+w_{1}w_{3}+w_{2}w_{4}+w_{1}w_{4}$. Increase $k$ by $1$.
(5-3) If it forms an induced cycle of length 1 mod 3 with the previously assigned $X$-3-path in $G_{k}$, then proceed as follows. Let $D_{1}$ be the induced cycle of length 1 mod 3. If $|V(P-D_{1})|=1$, then set $u\in V(P-D_{1})$ and add all edges $uv$ such that $v\in V(D_{1}-P)$ to $G_{k}$. Otherwise, take a 4-path $Q\subseteq D_{1}$ such that $|V(P\cap Q)|\geq 3$, and set $Q=w_{1}w_{2}w_{3}w_{4}$. Add two edges $w_{1}w_{3}$ and $w_{2}w_{4}$ to $G_{k}$. Assign $X$-3-paths to the induced cycles connecting without seams in $D_{1}+w_{1}w_{3}+w_{2}w_{4}$. Increase $k$ by $1$.
(6) Repeat (5) while they occur.
(7) Return the resulting graph $G_{k}$.
Let ${\bf K}(G)$ be a graph constructed by applying ${\bf K}$ to $G$. Note that ${\bf K}(G)$ is not unique and constructed from $G$ arbitrarily. By the scheme, ${\bf K}(G)$ is 2-connected with $V(G)=V({\bf K}(G))$ and $E(G)\subseteq E({\bf K}(G))$, and all edges in ${\bf K}(G)$ are covered by at least one set of $X$-3-paths but may be covered by other set of $X$-3-paths.
Remark 3.1.
Let $X$ be a dominating set of $G$. Every subset $D\subseteq X$ is a d-set of $G[N_{G}[D]]$ if and only if $X$ is a d-set of $G$.
Consider labeling vertices by assigning $X$-3-paths to the induced cycles of length 0 mod 3 connecting without seams in ${\bf K}(G)$ (Figure 1). Note that certain induced cycles are not counted for the labeling, and have as few labeled vertices as possible. The labeling case indicates the size of the set of all labeled vertices.
Figure 1: Labeling vertices on 0 mod 3 cycles connecting without seams
Lemma 3.1.
There exist at most $|V(G)|$ cases of labeling. (i) For every labeling, the set of all labeled vertices is a minimal dominating set of ${\bf K}(G)$. In addition, by selecting the first vertex to label, the number of the remaining labeled vertices are uniquely determined. (ii) For at least one labeling, the set of all labeled vertices is a d-set of ${\bf K}(G)$.
Proof.
Since every edge in ${\bf K}(G)$ is covered by $X$-3-path, for every labeling, the set of all labeled vertices is an independent dominating set of ${\bf K}(G)$, that is, a minimal dominating set of ${\bf K}(G)$. After selecting the vertex to label first, if there are multiple ways to label vertices during the labeling process, those vertices are interchangeable and equivalent. Thus, the number of the remaining labeled vertices is uniquely determined. Therefore, the statement (i) holds and there exist at most $|V(G)|$ cases of labeling. By Remark 3.1, the statement (ii) obviously holds.
Let $Y$ be a d-set of ${\bf K}(G)$ that is obtained by labeling vertices. Let $\mathcal{Y}$ be the set of all $Y$.
Proposition 3.1.
$\mathcal{Y}$ is determined in polynomial time and $|\mathcal{Y}|\leq|V(G)|$.
Proof.
By the definition of ${\bf K}$, ${\bf K}(G)$ is constructed from $G$ in polynomial time. By Lemma 3.1, there exist at most $|V(G)|$ cases of labeling, thus $\mathcal{Y}$ is determined in polynomial time and $|\mathcal{Y}|\leq|V(G)|$. ∎
Let $X$ be a d-set of $G$. Let $\mathcal{X}$ be the set of all $X$.
Lemma 3.2.
For some $X\in\mathcal{X}$, and some $Y\in\mathcal{Y}$, $Y\subseteq X$.
Proof.
For some $Y\in\mathcal{Y}$, if $Y$ is a dominating set of $G$, then for some $X\in\mathcal{X}$, $X=Y$. Otherwise, for all $Y\in\mathcal{Y}$, $Y$ is not a dominating set of $G$. Now, $Y$ is a d-set of $G[N_{G}[Y]]$. Let $W$ be a subset of $V(G)\setminus Y$ with minimum size such that $Y\cup W$ is a dominating set of $G$. By Remark 3.1, $Y\cup W$ is a d-set of $G$. ∎
Suppose that $Y$ is the set obtained by Lemma 3.2.
Lemma 3.3.
Let $G^{\prime}$ be a graph obtained by deleting $Y$ from, and adding edges to all vertex pairs of $\bigcup_{y\in Y}N_{G}(y)$ to $G$. Let $Z_{1}$ be a d-set of $G^{\prime}$. Let $G^{\prime\prime}=G-N_{G}[Y]$ and $Z_{2}$ be a d-set of $G^{\prime\prime}$. If $|Z_{1}|<|Z_{2}|$, then $Y\cup Z_{1}$ is a d-set of $G$, and $Z_{1}\cap\bigcup_{y\in Y}N_{G}(y)\neq\emptyset$. If $|Z_{1}|\geq|Z_{2}|$, then $Y\cup Z_{2}$ is a d-set of $G$.
Proof.
Obviously, $Y\cup Z_{2}$ is a d-set of $G$ if $|Z_{1}|\geq|Z_{2}|$. For a set $A\subseteq V(G)$, suppose that $Y\cup A$ is a d-set of $G$. Let $A^{\prime}=A\setminus N_{G}[Y]$. Suppose that $|Z_{1}|<|Z_{2}|$. Now, $A^{\prime}$ is not a dominating set of $G^{\prime\prime}$, and $A\cap\bigcup_{y\in Y}N_{G}(y)\neq\emptyset$. By the definition of $G^{\prime}$, it suffices that $A=Z_{1}$. ∎
Suppose that $G$ is cubic.
Theorem 3.1.
The d-set of $G$ is determined in polynomial time.
Proof.
By Lemma 3.2, for some $X\in\mathcal{X}$, and some $Y\in\mathcal{Y}$, $Y\subseteq X$. Let $G_{0}=G$. Let $G_{1}$ be constructed by deleting $Y$, and for every pair $w_{1},w_{2}\in\bigcup_{y\in Y}N_{G}(y)$, adding an edge $w_{1}w_{2}$ to $G_{0}$. Let $G_{2}=G_{0}-N_{G_{0}}[Y]$. Let $W_{1}$ be a d-set of $G_{1}$ and $W_{2}$ be a d-set of $G_{2}$. Since $G_{0}$ is cubic and by the definition of $G_{2}$, each component of $G_{2}$ is a path, cycle, or tree, or $G_{2}=\emptyset$. Thus $W_{2}$ is determined in polynomial time. Suppose that $|W_{1}|<|W_{2}|$. Let $Y_{1}$ be a d-set of ${\bf K}(G_{1})$ that satisfies Lemma 3.2. By the definition of $Y_{1}$, it suffices that $Y_{1}\cap\bigcup_{y\in Y}N_{G_{0}}(y)\neq\emptyset$. Let $G_{3}$ be constructed by deleting $Y_{1}$, and for every pair $w_{1},w_{2}\in\bigcup_{y\in Y_{1}}N_{G_{1}}(y)$, adding an edge $w_{1}w_{2}$ to $G_{1}$. Let $G_{4}=G_{1}-N_{G_{1}}[Y_{1}]$. Let $W_{3}$ be a d-set of $G_{3}$ and $W_{4}$ be a d-set of $G_{4}$. By the definition of $G_{4}$, each component of $G_{4}$ is a path, or $G_{4}=\emptyset$. Thus $W_{4}$ is determined in polynomial time. Suppose that $|W_{3}|<|W_{4}|$. Let $Y_{2}$ be a d-set of ${\bf K}(G_{3})$ that satisfies Lemma 3.2. By the definition of $Y_{2}$, it suffices that $Y_{2}\cap\bigcup_{y\in Y_{1}}N_{G_{1}}(y)\neq\emptyset$. Let $G_{5}$ be constructed by deleting $Y_{2}$, and for every pair $w_{1},w_{2}\in\bigcup_{y\in Y_{2}}N_{G_{3}}(y)$, adding an edge $w_{1}w_{2}$ to $G_{3}$. Let $G_{6}=G_{3}-N_{G_{3}}[Y_{2}]$. Let $W_{5}$ be a d-set of $G_{5}$ and $W_{6}$ be a d-set of $G_{6}$. By the definition of $G_{6}$, $G_{6}$ is independent, or $G_{6}=\emptyset$. Thus $W_{6}=V(G_{6})$. Suppose that $|W_{5}|<|W_{6}|$. Let $Y_{3}$ be a d-set of ${\bf K}(G_{5})$ that satisfies Lemma 3.2. By the definition of $Y_{3}$, it suffices that $Y_{3}\cap\bigcup_{y\in Y_{2}}N_{G_{3}}(y)\neq\emptyset$. Now, $Y_{3}$ is a dominating set of $G_{5}$ and so it suffices that $W_{5}=Y_{3}$. By Lemma 3.3, if $|W_{5}|<|W_{6}|$, then it suffices that $W_{3}=Y_{2}\cup W_{5}$, otherwise, it suffices that $W_{3}=Y_{2}\cup W_{6}$. By Lemma 3.3, if $|W_{3}|<|W_{4}|$, then it suffices that $W_{1}=Y_{1}\cup W_{3}$, otherwise, it suffices that $W_{1}=Y_{1}\cup W_{4}$. By Lemma 3.3, if $|W_{1}|<|W_{2}|$, then it suffices that $X=Y\cup W_{1}$, otherwise, it suffices that $X=Y\cup W_{2}$. By Proposition 3.1, $Y$ is determined in polynomial time. Thus, $X$ is determined in polynomial time. The proof is completed. ∎